The restaurant at the top of a very tall building in Seattle rotates 0.99 revolutions every hour. Miranda sits 20.8 meters from the center of the restaurant near a window.
(a1) Through how many radians does Miranda turn in 108 minutes?
(exact) angle = ?rads
(a2) Now, report the answer accurate to 3 decimal places:
(approximate) angle = ?rads
(b1) How far does Miranda move in 108 minutes?
(exact) distance = ? meters
(b2) Now, report the answer accurate to 1 decimal place:
(approximate) distacne = ?meters

Answer by rothauserc(4718)

1 revolution is 360 degrees
:
(a1) 360 * 0.99 / 60 = 5.94 degrees / minute
:
108 * 5.94 = 641.52 degrees
:
641.52 = 641.52 * pi / 180 = 11.1966 radians
:
(a2) 641.52 = 11.197 radians
:
(b1) distance Miranda travels in 1 revolution = 2 * pi * 20.8 = 130.6903 meters
:
(641.52 / 360) * 130.6903 = 232.8901 meters
:
(b2) 232.9 meters
: